OpenStudy (anonymous):

Help me??? Given f (x) = 3x^3 - 4x^2 + x-1 what is f(2) Factor 36n^2 - 81 Simplify : mm(m^5)^0 m^-3 divided by m^6(y^3)^-3 solve : y^2 = 18

OpenStudy (campbell_st):

f(2) is simply a case of substituting x = 2 into the equation and evaluating. 36n^2 - 81 this is the difference of 2 squares 36n^2 = (6n)^2 81 = (9)^2 hope it helps

OpenStudy (anonymous):

for my first question my answer was 3, am I right?

OpenStudy (campbell_st):

I don't think so \[3 \times2^3 - 4 \times 2^2 + 2 - 1\]